What volume of oxygen is required for the complete combustion of mixture of 5cm3 of CH4 and 5cm3 of C2H4
1
Expert's answer
2012-10-22T12:21:17-0400
v1 & 5
2O2 + CH4 = CO2 + 2H2O

& v2 5
5/2O2 + C2H4 = 2CO2 + H2O

v1=5*2=10 cm3

v2=5* 5/2=12.5

v=v1+v2=22.5 cm3